OpenStudy (anonymous):

Transform Each Function into the form f(x)=a(x-h)^2+k 1. f(x)=x^2-2x-1 2. f(x)=x^2-6x+4 3. f(x)=x^2+4x+2 4. f(x)=3x^2-5x+1 5. f(x)=2x^2-5x+3 6. f(x)=2x^2-3x-1

OpenStudy (anonymous):

1. (x-1)^2 - 1 - 1 = (x-1)^2 - 2 this is how it works: divide the coefficient of x in f(x) - in this case it is -2 - by 2 to give -1 - now we wrte (x-1)^2 which when expanded gives +1 as last term - so we need to subtract 1 to cancel this finally we add -1 so that it is identical to original f(x)

OpenStudy (anonymous):

work out 2 and 3 in similar way

OpenStudy (anonymous):

i don't get it pls explain further or just pls answer all so i can get it we have exam tommorow i have to review this please :)

OpenStudy (anonymous):

ok - first one is x^2 - 2x -1 divide the -2 in middle by 2 this gives -1 OK?

OpenStudy (anonymous):

ok . i get it know .

OpenStudy (anonymous):

so we write (x-1)^2 - if we expand this it gives x^2 - 2x + 1 - so to get back to x^2 -2x we have to subtract 1.

OpenStudy (anonymous):

(x-1)^2 - 1 is the same as x^2 - 2x only written in a different form

OpenStudy (anonymous):

now to get back to the original f(x) we have to take another 1 away so the final resukt is (x-1)^2 - 1 - 1 = (x-1)^2 - 2

OpenStudy (anonymous):

illl start the second one for you: f(x)=x^2-6x+4 = (x - 3)^2 (-3 because its -6 / 2)

OpenStudy (anonymous):

and now you subtract 16 and add 4